WEB(1) WEB(1) NAME tangle,weave−translateWEBtoPascaland/orTeX SYNOPSIS tanglewebfile[.web][changefile[.ch]] weave[-x]webfile[.web][changefile[.ch]] DESCRIPTION ThetangleprogramconvertsaWEBsourcedocumentintoaPascalprogramthatmaybecompiledintheusualwaywiththeon-linePascalcompiler(e.g.,pc(1)).Theoutputfileisallin lowercaseandpackedintolinesof72charactersorless,withtheonlyconcessiontoreadability beingtheterminationoflinesatsemicolonswhenthiscanbedoneconveniently. TheWEBlanguageallowsyoutoprepareasingledocumentcontainingalltheinformationthatis neededbothtoproduceacompilablePascalprogramandtoproduceawell-formatteddocument describingtheprograminasmuchdetailasthewritermaydesire.TheuserofWEBmustbe familiarwithbothTeXandPascal.WEBalsoprovidesarelativelysimple,althoughadequate, macrofacilitythatpermitsaPascalprogramtobewritteninsmalleasily-understoodmodules. Thecommandlineshouldhaveeitheroneortwonamesonit.ThefirstistakenastheWEBfile (and.webisaddedifthereisnoextension).Ifthereisanothername,itisachangefile(and.ch isaddedifthereisnoextension).ThechangefileoverridespartsoftheWEBfile,asdescribedin theWEBsystemdocumentation. TheoutputfilesareaPascalfileandastringpoolfile,whosenamesareformedbyadding.pand .poolrespectivelytotherootoftheWEBfilename. TheweaveprogramisusedtocreateaTeXfileforviewingtheWEBprogram.Ittakesappropriatecareoftypographicdetailslikepagelayoutandtheuseofindentation,italics,boldface,etc., anditsuppliesextensivecross-indexinformationthatitgathersautomatically.Thecommandline argumentsarethesameasfortangleexceptfortheoption:−xsaystoomittheindex,module namelist,andtableofcontentspages.(ACONTENTS.texfilewillstillbewrittenwhentheTeX fileisprocessed,however,unlesssomemacrosinwebmac.texareredefined.) TheoutputTeXfilenameisformedbyadding.textotherootoftheWEBfilename. Thereareseveralmacrosthatprobablyshouldberedefinedbytheprogrammeratthebeginning oftheWEBfile.Itisagoodideatoset\titletothenameoftheprogram.And,tocauseoutput ofonlychangedmodules,onecansay\let\maybe=\iffalse(usuallyasthefirstchangeinthe changefile). FILES /usr/TeX/lib/tex/macros/webmac.texTeXmacrosusedbyweaveoutput. SEEALSO TheWEBSystemofStructuredDocumentationandLiterateProgramming,byD.E.Knuth. WeavingaProgram,byWayneSewell tex(1),pc(1) pxp(1)(forformattingtangleoutputwhendebugging) TeX:TheProgramandMETAFONT:TheProgram,volumesBandDintheComputersandTypesettingseries,publishedbyAddison-Wesley,arebyfarthelargestextantexamplesofWEBpro-grams. AUTHORS WEBwasdesignedbyDonaldE.Knuth,basedonanearliersystemcalledDOC(implementedby IgnacioZabala).ThetangleandweaveprogramsarethemselveswritteninWEB.Thesystemwas originallyportedtoUnixatStanfordbyHowardTrickey,andatCornellbyPavelCurtis. 7Jan92 1